MS. SIM PAID 55 DOLLARS FOR A SKIRT AND 2 BLOUSES. THE SKIRT COST 5.50 DOLLARS MORE THAN EACH BLOUSE. FIND THE COST OF THE BLOUSE
Answer by mananth(16949)

Blouse = x
Skirt = x+5.5
x+5.5+2x=55
3x+5.5=55
3x=49.5
x=49.5/3
x= 16.50
Blouse costs $ 16.50
m.ananth@hotmail.ca